OpenStudy (anonymous):

What is the common ratio of the sequence: 2, -6, 18, -54? why? and what are the next 3 terms in the sequence?

jimthompson5910 (jim_thompson5910):

To find the common ratio, take any term and divide it by the previous term

OpenStudy (anonymous):

what @jim_thompson5910 said so you can use \(\frac{-6}{2}\) or \(\frac{18}{-6}\) or \(\frac{-54}{18}\) as they are all equal

OpenStudy (anonymous):

thank you guys :)

jimthompson5910 (jim_thompson5910):

yw